# Global Bifurcation of Non-Radial Solutions for Symmetric Sub-linear Elliptic Systems on the Planar Unit Disc

### v2

@@ 28 View(max_orbtyps);
For any $m>0$ and $j\in\{0,2,3\}$, the output of the above program can be used to describe the the set of maximal orbit types $\mathfrak{M}_{m,j}\subset\Phi_{0}(G)$ the basic degrees $\deg_{\mathcal{V}_{m,j}}\in A(G)$ (cf. (24) and (25)), as follows:
@@
We can deduce, from the coefficients of the Basic Degrees $\deg_{\mathcal{V}_{1,j}}\in A(G)$ (cf. Remark 4.3), that the fixed point space $\mathcal{V}_{1,j}^{H}$ has odd dimension for any $j\in\{0,2,3\}$ and $(H)\in\mathfrak{M}_{1,j}$. Since the critical set (51) is manageably small, we can manually compute the cardinalities (27) for each of the critical points $(\alpha_{n,m,j},0)\in\Lambda$ to obtain the quantities
$\mathfrak{s}(\alpha_{1,3,2},H)$ $=\mathfrak{s}(\alpha_{1,3,3},H)=\mathfrak{s}(\alpha_{1,3,0},H)=3,$
$\mathfrak{s}(\alpha_{2,1,2},H)$ $=\mathfrak{s}(\alpha_{2,1,3},H)=1.$
@@
$\mathfrak{i}^{3}(\alpha_{1,3,2},H)$ $=1,\quad\mathfrak{i}^{3}(\alpha_{1,3,3},H)=-1,\quad\mathfrak{i}^{3}(\alpha_{1,3,0},H)=1,$
$\mathfrak{i}^{1}(\alpha_{2,1,2},H)$ $=1,\quad\mathfrak{i}^{1}(\alpha_{2,1,3},H)=-1.$
Therefore, from Theorem 4.1, it follows that every critical point $(\alpha_{n,m,j},0)\in\Lambda$ is a branching point for branches of non-trivial solutions to the problem (42) with symmetries at least $({}^{3}H)$ in the case of $(\alpha_{1,3,2},0),(\alpha_{1,3,3},0),(\alpha_{1,3,0},0)\in\Lambda$ and with symmetries at least $(H)$ in the case of $(\alpha_{2,1,2},0),(\alpha_{2,1,3},0)\in\Lambda$. The following G.A.P code can be used to computationally verify the non-triviality of the local bifurcation invariant (specifically, the non-triviality of the relevant coefficients) at each of the critical points in terms of the basic degrees
@@
2# Initialize the Burnside Ring A(G)
3AG:=BurnsideRing(G); H:=Basis(AG);
4# Initialize the unit element (G) $\in$ A(G)
5U:=H[0,131];
6# Initialize the relevant Basic Degrees
7d_30 := BasicDegree(Irr(G)[3,3]);
8d_32 := BasicDegree(Irr(G)[3,5]);
9d_33 := BasicDegree(Irr(G)[3,7]);
10d_12 := BasicDegree(Irr(G)[1,5]);
11d_13 := BasicDegree(Irr(G)[1,7]);
12# Compute Products of Basic Degrees Cumulatively
13p_1 := d_32;
14p_2 := p_1*d_33;
15p_3 := p_2*d_30;
16p_4 := p_3*d_12;
17p_5 := p_4*d_13;
@@
19w_132 := U - p_1;
20w_133 := p_1 - p_2;
21w_130 := p_2 - p_3;
22w_212 := p_3 - p_4;
23w_213 := p_4 - p_5;
24# Sum the local bifurcation invariants to determine the Rabinowitz Alternative
25sum := w_132 + w_133 + w_130 + w_212 + w_213
@@
$\omega_{G}(\alpha_{1,3,2})$ $=-4(D_{6}^{{\mathbb{Z}}_{3}}\times^{V_{4}}D_{4}^{p})-(D_{6}^{D_{3}}\times^{V_{4}}V_{4}^{p})+(D_{6}^{D_{3}}\times^{D_{4}}D_{4}^{p})+(D_{6}^{D_{3}}\times^{D_{4}^{\hat{d}}}D_{4}^{p})+2(D_{18}^{{\mathbb{Z}}_{3}}\times^{V_{4}}S_{4}^{p}),$
$\omega_{G}(\alpha_{1,3,3})$ $=4(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{1}}D_{1}^{p})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{1}}D_{2}^{p})+(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{1}}{\mathbb{Z}}_{1}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{D_{1}}D_{4}^{z})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{z})$
$\quad-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}^{-}}V_{4}^{p})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}{\mathbb{Z}}_{4}^{p})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}D_{4}^{z})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}V_{4}^{p})-(D_{6}^{D_{3}}\times^{D_{1}}D_{1}^{p})$
$\quad-(D_{6}^{D_{3}}\times^{D_{1}^{z}}D_{1}^{p})-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{2}^{-}}D_{2}^{p})+(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{2}}D_{2}^{p})-2(D_{18}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{1}}D_{3}^{p})+2(D_{12}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{p})$
$\quad-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{D_{2}^{z}}D_{4}^{p})+(D_{6}^{D_{3}}\times^{D_{2}^{d}}D_{4}^{z})-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})-(D_{6}^{D_{3}}\times^{D_{2}^{z}}D_{4}^{z})$
@@
$\omega_{G}(\alpha_{1,3,0})$ $=2(D_{6}^{{\mathbb{Z}}_{3}}\times^{D_{1}}D_{4}^{z})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}{\mathbb{Z}}_{4}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}D_{4}^{z})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}V_{4}^{p})+(D_{6}^{D_{3}}\times^{D_{1}}D_{1}^{p})$
$\quad-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{2}}D_{2}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{3}}D_{3}^{p})-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{3}}{\mathbb{Z}}_{3}^{p})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})+3(D_{6}^{{\mathbb{Z}}_{3}}\times^{V_{4}}D_{4}^{p})$
$\quad+(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})+(D_{6}^{D_{3}}\times^{V_{4}}V_{4}^{p})-2(D_{6}^{D_{3}}\times^{D_{4}}D_{4}^{p})+(D_{6}^{D_{3}}\times^{S_{4}}S_{4}^{p})$
$\omega_{G}(\alpha_{2,1,2})$ $=2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{2}^{p})+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{1}}{\mathbb{Z}}_{1}^{p})-2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}}D_{4}^{z})-2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}^{z}}D_{4}^{z})-(D_{2}^{D_{1}}\times^{D_{1}}D_{1}^{p})$
$\quad-(D_{2}^{D_{1}}\times^{D_{1}^{z}}D_{1}^{p})+4(D_{6}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{3}^{p})-2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{2}^{z}}D_{4}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{V_{4}}D_{4}^{p})$
$\quad-(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})+(D_{2}^{D_{1}}\times^{D_{2}^{z}}D_{4}^{z})+(D_{2}^{D_{1}}\times^{D_{4}}D_{4}^{p})-(D_{2}^{D_{1}}\times^{D_{4}^{\hat{d}}}D_{4}^{p})-2(D_{6}^{{\mathbb{Z}}_{1}}\times^{V_{4}}S_{4}^{p})$
$\omega_{G}(\alpha_{2,1,3})$ $=-4(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{1}^{p})-4(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{2}^{p})-2(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{1}}{\mathbb{Z}}_{1}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}}D_{4}^{z})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}^{z}}D_{4}^{z})$
$\quad+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{z})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{2}^{-}}V_{4}^{p})+(D_{2}^{D_{1}}\times^{D_{1}}D_{1}^{p})+2(D_{2}^{D_{1}}\times^{D_{1}^{z}}D_{1}^{p})+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{2}^{-}}D_{2}^{p})$
$\quad-2(D_{6}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{3}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{3}}D_{3}^{p})+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{3}}{\mathbb{Z}}_{3}^{p})-2(D_{4}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})$
$\quad-(D_{2}^{D_{1}}\times^{D_{2}^{d}}D_{4}^{z})+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})+(D_{2}^{D_{1}}\times^{D_{3}^{z}}D_{3}^{p})-(D_{2}^{D_{1}}\times^{D_{4}^{z}}D_{4}^{p}).$
@@
Next, we can determine the global behaviour of some of the branches predicted using Theorem (4.1) by establishing the parity of the set (39). Notice that $2\nmid\max\limits_{(\alpha,0)\in\Lambda}\{\mathfrak{s}(\alpha,H)\}$ is obtained for exactly three critical points $(\alpha_{1,3,0},0),(\alpha_{1,3,2},0),(\alpha_{1,3,3},0)$, i.e.
$\bar{\mathfrak{s}}(H)=3\;\text{ and }\;\mathfrak{J}(H)=\{(\alpha_{1,3,0},0),(\alpha_{1,3,2},0),(\alpha_{1,3,3},0)\}.$
@@
Therefore, any branch of non-trivial solutions with symmetries at least $({}^{3}H)$ bifurcating from $(\alpha_{1,3,0},0),(\alpha_{1,3,2},0),(\alpha_{1,3,3},0)$ consists only of non-radial solutions and is unbounded. Of course, since the critical set (51) is finite, one needs only sum the relevant local bifurcation invariants to arrive at the same conclusion directly from the Rabinowitz alternative (cf. Theorems 4.2 and 4.3).
@@
In conclusion, let’s consider an interpretation of the above results. Take the maximal orbit type $(H)=(D_{2}^{D_{1}}\times^{D_{4}}D_{4}^{p})\in\mathscr{M}_{1,2}$ and let $\mathcal{C}$ be a branch of non-radial solutions with symmetries at least $({}^{3}H)=(D_{6}^{D_{3}}\times^{D_{4}}D_{4}^{p})\in\mathscr{M}_{3,2}$ emerging from the first critical point $(\alpha_{1,3,2},0)\in\Lambda$. The kernel of $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ at any critical point $(\alpha_{n,m,j},0)\in\Lambda$ is given by the corresponding irreducible $G$-representation $\mathscr{E}_{n,m}^{j}$ (cf. (15)), i.e.
Ker $\text{\rm Ker\,}\mathscr{A}(\alpha_{1,3,2})=\mathscr{E}_{1,3}^{2}.$
In particular, any vector $\hat{u}\in\mathscr{H}$ with $\mathscr{A}(\alpha_{1,3,2})\hat{u}=0$ must be of the form
@@
where $\vec{w}_{1},\vec{w}_{2}\in\mathcal{V}_{2}^{-}$ are the eigenvectors $\vec{w}_{1}=(-1,1,-1,1,0,0)^{T}$ and $\vec{w}_{2}=(1,0,1,0,-1,-1)^{T}$. Computing the generators of $(H)$ and examining their action on $\mathscr{E}_{1,3}^{2}$, it can be shown that $\hat{u}\in\text{\rm Ker\,}\mathscr{A}(\alpha_{0})^{H}$ if and only if $\vec{b}=0$ and $\vec{a}=\vec{w}_{1}+2\vec{w}_{2}$. In particular, one has $\dim\text{\rm Ker\,}\mathscr{A}(\alpha_{0})^{H}=1$, which implies, from the celebrated theorem of Crandall-Rabinowitz (cf. [29]), that the branch $\mathcal{C}$ is tangent to the solution $\hat{u}(r,\theta)=J_{3}(\sqrt{s_{13}}r)\cos(3\theta)(\vec{w}_{1}+2\vec{w}_{2})$ at the point $(\alpha_{1,3,2},0)$. Therefore, $\hat{u}\in\mathscr{H}$ can reasonably be used to estimate the behaviour of the branch $\mathcal{C}$ for parameter values near $\alpha_{1,3,2}\in\mathbb{R}$.

### v3

@@ 27 View(max_orbtyps);
For any $m>0$ and $j\in\{0,2,3\}$, the output of the above program can be used to describe the set of maximal orbit types $\mathfrak{M}_{m,j}\subset\Phi_{0}(G)$ the basic degrees $\deg_{\mathcal{V}_{m,j}}\in A(G)$ (cf. (25) and (26)), using amalgamated notation (cf. Appendix A), as follows:
@@
We can deduce, from the coefficients of the Basic Degrees $\deg_{\mathcal{V}_{1,j}}\in A(G)$ (cf. Remark 4.3), that the fixed point space $\mathcal{V}_{1,j}^{H}$ has odd dimension for any $j\in\{0,2,3\}$ and $(H)\in\mathfrak{M}_{1,j}$. Since the critical set (52) is manageably small, we also can manually compute the cardinalities (28) for each of the critical points $(\alpha_{n,m,j},0)\in\Lambda$ to obtain the quantities
$\begin{cases}\mathfrak{s}(\alpha_{1,3,0},H)=3\quad&\text{ if }(H)\in\mathfrak{M}_{1,0};\\ \mathfrak{s}(\alpha_{1,3,2},H)=3,\;\mathfrak{s}(\alpha_{2,1,2},H)=1\quad&\text{ if }(H)\in\mathfrak{M}_{1,2};\\ \mathfrak{s}(\alpha_{1,3,3},H)=3,\;\mathfrak{s}(\alpha_{2,1,3},H)=1\quad&\text{ if }(H)\in\mathfrak{M}_{1,3},\end{cases}$
@@
$\begin{cases}\mathfrak{i}^{3}(\alpha_{1,3,0},H)=1\quad&\text{ if }(H)\in\mathfrak{M}_{1,0};\\ \mathfrak{i}^{3}(\alpha_{1,3,2},H)=i^{1}(\alpha_{2,1,2},H)=1\quad&\text{ if }(H)\in\mathfrak{M}_{1,2};\\ \mathfrak{i}^{3}(\alpha_{1,3,3},H)=i^{1}(\alpha_{2,1,3},H)=-1\quad&\text{ if }(H)\in\mathfrak{M}_{1,3}.\end{cases}$
Therefore, from Theorem 4.1, it follows that every critical point $(\alpha_{n,m,j},0)\in\Lambda$ is a branching point for branches of non-trivial solutions to the problem (43) with symmetries at least $({}^{3}H)$ in the case of $(\alpha_{1,3,2},0),(\alpha_{1,3,3},0),(\alpha_{1,3,0},0)\in\Lambda$ and with symmetries at least $(H)$ in the case of $(\alpha_{2,1,2},0),(\alpha_{2,1,3},0)\in\Lambda$ for each orbit type of maximal kind $(H)\in\mathfrak{M}_{1,j}$. The following G.A.P code can be used to computationally verify the non-triviality of the local bifurcation invariant (specifically, the non-triviality of the relevant coefficients) at each of the critical points in terms of the basic degrees
@@
2# Initialize the Burnside Ring A(G), with unit element U =(G)
3AG:=BurnsideRing(G); H:=Basis(AG); U:=H[0,131];
4# Initialize the relevant Basic Degrees
5d_30 := BasicDegree(Irr(G)[3,3]);
6d_32 := BasicDegree(Irr(G)[3,5]);
7d_33 := BasicDegree(Irr(G)[3,7]);
8d_12 := BasicDegree(Irr(G)[1,5]);
9d_13 := BasicDegree(Irr(G)[1,7]);
10# Compute Products of Basic Degrees Cumulatively
11p_1 := d_32; p_2 := p_1*d_33; p_3 := p_2*d_30;
12p_4 := p_3*d_12; p_5 := p_4*d_13;
@@
14w_132 := U - p_1; w_133 := p_1 - p_2; w_130 := p_2 - p_3;
15w_212 := p_3 - p_4; w_213 := p_4 - p_5;
16# Sum the local bifurcation invariants
17sum := w_132 + w_133 + w_130 + w_212 + w_213
@@
$\omega_{G}(\alpha_{1,3,2})$ $=-4(D_{6}^{{\mathbb{Z}}_{3}}\times^{V_{4}}D_{4}^{p})-(D_{6}^{D_{3}}\times^{V_{4}}V_{4}^{p})+(D_{6}^{D_{3}}\times^{D_{4}}D_{4}^{p})+(D_{6}^{D_{3}}\times^{D_{4}^{\hat{d}}}D_{4}^{p})$
$\quad+2(D_{18}^{{\mathbb{Z}}_{3}}\times^{V_{4}}S_{4}^{p}),$
$\omega_{G}(\alpha_{1,3,3})$ $=4(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{1}}D_{1}^{p})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{1}}D_{2}^{p})+(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{1}}{\mathbb{Z}}_{1}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{D_{1}}D_{4}^{z})$
$\quad-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{z})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}^{-}}V_{4}^{p})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}{\mathbb{Z}}_{4}^{p})+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}D_{4}^{z})$
$\quad+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}V_{4}^{p})-(D_{6}^{D_{3}}\times^{D_{1}}D_{1}^{p})-(D_{6}^{D_{3}}\times^{D_{1}^{z}}D_{1}^{p})-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{2}^{-}}D_{2}^{p})$
$\quad+(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{2}}D_{2}^{p})-2(D_{18}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{1}}D_{3}^{p})+2(D_{12}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})$
$\quad-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{D_{2}^{z}}D_{4}^{p})+(D_{6}^{D_{3}}\times^{D_{2}^{d}}D_{4}^{z})-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})-(D_{6}^{D_{3}}\times^{D_{2}^{z}}D_{4}^{z})$
@@
$\omega_{G}(\alpha_{1,3,0})$ $=2(D_{6}^{{\mathbb{Z}}_{3}}\times^{D_{1}}D_{4}^{z})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}{\mathbb{Z}}_{4}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}D_{4}^{z})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{2}}V_{4}^{p})$
$\quad+(D_{6}^{D_{3}}\times^{D_{1}}D_{1}^{p})-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{2}}D_{2}^{p})-2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{3}}D_{3}^{p})-(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{3}}{\mathbb{Z}}_{3}^{p})$
$\quad+2(D_{6}^{{\mathbb{Z}}_{3}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})+3(D_{6}^{{\mathbb{Z}}_{3}}\times^{V_{4}}D_{4}^{p})+(D_{6}^{D_{3}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})+(D_{6}^{D_{3}}\times^{V_{4}}V_{4}^{p})$
$\quad-2(D_{6}^{D_{3}}\times^{D_{4}}D_{4}^{p})+(D_{6}^{D_{3}}\times^{S_{4}}S_{4}^{p})$
$\omega_{G}(\alpha_{2,1,2})$ $=2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{2}^{p})+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{1}}{\mathbb{Z}}_{1}^{p})-2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}}D_{4}^{z})-2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}^{z}}D_{4}^{z})$
$\quad-(D_{2}^{D_{1}}\times^{D_{1}}D_{1}^{p})-(D_{2}^{D_{1}}\times^{D_{1}^{z}}D_{1}^{p})+4(D_{6}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{3}^{p})-2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})$
$\quad+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{2}^{z}}D_{4}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{V_{4}}D_{4}^{p})-(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})+(D_{2}^{D_{1}}\times^{D_{2}^{z}}D_{4}^{z})$
$\quad+(D_{2}^{D_{1}}\times^{D_{4}}D_{4}^{p})-(D_{2}^{D_{1}}\times^{D_{4}^{\hat{d}}}D_{4}^{p})-2(D_{6}^{{\mathbb{Z}}_{1}}\times^{V_{4}}S_{4}^{p})$
$\omega_{G}(\alpha_{2,1,3})$ $=-4(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{1}^{p})-4(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{2}^{p})-2(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{1}}{\mathbb{Z}}_{1}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}}D_{4}^{z})$
$\quad+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{D_{1}^{z}}D_{4}^{z})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{z})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{2}^{-}}V_{4}^{p})+(D_{2}^{D_{1}}\times^{D_{1}}D_{1}^{p})$
$\quad+2(D_{2}^{D_{1}}\times^{D_{1}^{z}}D_{1}^{p})+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{2}^{-}}D_{2}^{p})-2(D_{6}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{1}}D_{3}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{3}}D_{3}^{p})$
$\quad+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{3}}{\mathbb{Z}}_{3}^{p})-2(D_{4}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{2}^{-}}D_{4}^{p})+2(D_{2}^{{\mathbb{Z}}_{1}}\times^{{\mathbb{Z}}_{4}}D_{4}^{p})-(D_{2}^{D_{1}}\times^{D_{2}^{d}}D_{4}^{z})$
$\quad+(D_{2}^{D_{1}}\times^{{\mathbb{Z}}_{4}}{\mathbb{Z}}_{4}^{p})+(D_{2}^{D_{1}}\times^{D_{3}^{z}}D_{3}^{p})-(D_{2}^{D_{1}}\times^{D_{4}^{z}}D_{4}^{p}).$
@@
Next, we can determine the global behaviour of some of the branches predicted using Theorem (4.1) by establishing the parity of the set (40). Notice that, for every $j\in\{0,2,3\}$ and $(H)\in\mathfrak{M}_{1,j}$, the quantity $\bar{\mathfrak{s}}(H)=3$ is obtained for exactly one critical point $(\alpha_{1,3,j},0)$, i.e.
$\mathfrak{J}(H)=\begin{cases}\{(\alpha_{1,3,0},0)\}\quad&\text{ if }(H)\in\mathfrak{M}_{1,0};\\ \{(\alpha_{1,3,2},0)\}\quad&\text{ if }(H)\in\mathfrak{M}_{1,2};\\ \{(\alpha_{1,3,3},0)\}\quad&\text{ if }(H)\in\mathfrak{M}_{1,3}.\end{cases}$
@@
Therefore, any branch of non-trivial solutions with symmetries at least $({}^{3}H)$ bifurcating from $(\alpha_{1,3,j},0)$, for every $j\in\{0,2,3\}$ and $(H)\in\mathfrak{M}_{1,j}$, consists only of non-radial solutions and is unbounded. Of course, since the critical set (52) is finite, one needs only sum the relevant local bifurcation invariants to arrive at the same conclusion directly from the Rabinowitz alternative (cf. Theorems 4.2 and 4.3).
@@
In closing, let’s consider a closer interpretation of the above results. Take the maximal orbit type $(H)=(D_{2}^{D_{1}}\times^{D_{4}}D_{4}^{p})\in\mathfrak{M}_{1,2}$ and let $\mathcal{C}$ be a branch of non-radial solutions with symmetries at least $({}^{3}H)=(D_{6}^{D_{3}}\times^{D_{4}}D_{4}^{p})\in\mathfrak{M}_{3,2}$ emerging from the first critical point $(\alpha_{1,3,2},0)\in\Lambda$. Notice that the kernel of $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ at any critical point $(\alpha_{n,m,j},0)\in\Lambda$ is given by the corresponding irreducible $G$-representation $\mathscr{E}_{n,m}^{j}$ (cf. (16)), i.e.
Ker $\text{\rm Ker\,}\mathscr{A}(\alpha_{1,3,2})=\mathscr{E}_{1,3}^{2}$
such that any vector $\hat{u}\in\mathscr{H}$ with $\mathscr{A}(\alpha_{1,3,2})\hat{u}=0$ must be of the form
@@
where $\vec{w}_{1},\vec{w}_{2}\in\mathcal{V}_{2}^{-}$ are the eigenvectors
$\vec{w}_{1}=(-1,1,-1,1,0,0)^{T}\text{ and }\vec{w}_{2}=(1,0,1,0,-1,-1)^{T}.$
Computing the generators of the orbit type $(D_{6}^{D_{3}}\times^{D_{4}}D_{4}^{p})$,
$\left(\frac{2\pi}{3},(1),1\right),\;\left(\kappa,(1),1\right),\;\left(e_{SO(2)},(1,2,3,4),1\right),\;\left(e_{SO(2)},(1,2),1\right),\;\left(\frac{\pi}{3},(1),-1\right)\in H$
and examining their action (cf. (14), (45)) on $\mathscr{E}_{1,3}^{2}$, we observe the radial invariances
$\hat{u}(r,\theta+\frac{2\pi}{3})=\hat{u}(r,\theta)\text{ and}-\hat{u}(r,\theta+\frac{\pi}{3})=\hat{u}(r,\theta),$
which are satisfied trivially for all elements of $\mathscr{E}_{1,2}^{3}$, the conjugate invariance
$\hat{u}(r,-\theta)=\hat{u}(r,\theta),$
which is only satisfied when $\vec{b}=\vec{0}$, and also the permutation invariances
$(1,2,3,4)\cdot\hat{u}(r,\theta)=\hat{u}(r,\theta)\text{ and }(1,2)\cdot\hat{u}(r,\theta)=\hat{u}(r,\theta),$
which, for $\vec{a}=a_{1}\vec{w}_{1}+a_{2}\vec{w}_{2}$, is satisfied if and only if $a_{2}=2a_{1}$. Moreover, since one has $\dim\text{\rm Ker\,}\mathscr{A}(\alpha_{0})^{H}=1$, the celebrated theorem of Crandall-Rabinowitz (cf. [29]) guarantees that the branch $\mathcal{C}$ is tangent to the solution $\hat{u}(r,\theta)=J_{3}(\sqrt{s_{13}}r)\cos(3\theta)(\vec{w}_{1}+2\vec{w}_{2})$ at the point $(\alpha_{1,3,2},0)$. Therefore, $\hat{u}\in\mathscr{H}$ can reasonably be used to estimate the behaviour of the branch $\mathcal{C}$ for parameter values near $\alpha_{1,3,2}\in\mathbb{R}$.